# LiDAR Depth Map Guided Image Compression Model

### 12 Jan 2024 v1

@@ Abstract
The incorporation of LiDAR technology into some high-end smartphones has unlocked numerous possibilities across various applications, including photography, image restoration, augmented reality, and more. In this paper, we introduce a novel direction that harnesses LiDAR depth maps to enhance the compression of the corresponding RGB camera images. Specifically, we propose a Transformer-based learned image compression system capable of achieving variable-rate compression using a single model while utilizing the LiDAR depth map as supplementary information for both the encoding and decoding processes. Experimental results demonstrate that integrating LiDAR yields an average PSNR gain of $0.83$ dB and an average bitrate reduction of $16\%$ as compared to its absence.
Index Terms:
Learned image compression, LiDAR, depth map, prompts, transformer.
I Introduction
@@
Figure 1: High-level architectural framework. The image compression model takes input from both the RGB image and the LiDAR depth data, using the LiDAR map to enhance the compression process of the RGB image. In the considered scenario, the LiDAR map is available at both the encoder and the decoder.
@@
The rest of the paper is structured as follows. Sec. II depicts the specific scenario under investigation. Sec. III introduces the architecture proposed for incorporating LiDAR information, and Sec. IV shows the experimental results. Finally, Sec. V provides the concluding remarks.
II Examined Scenario
@@
Let us consider the simple scenario to verify whether the LiDAR map can truly assist in the compression of an RGB image. Thus, let us assume the LiDAR map is accessible at both the encoder and the decoder, as depicted in Fig. 1. Hence, we examine the situation where the RGB image and the corresponding LiDAR map are simultaneously acquired from the device, and the LiDAR data assists in both encoding the RGB image, which is then locally stored on the device, and decoding the reconstructed version. Note that these data are often acquired simultaneously due to the LiDAR map’s utility in various applications, as previously mentioned. In essence, our approach capitalizes on the presence of the LiDAR map on the device to enhance the compression of the RGB image.
@@
Figure 2: The network architecture of the proposed LiDAR depth map guided image compression model. For variable rate control, the prompt generation networks $p_{a}$ and $p_{s}$ produce prompt tokens for the encoder $g_{a}$ and decoder $g_{s}$, respectively. Two additional prompt generation networks, denoted as $l_{a}$ and $l_{s}$, receive the LiDAR map as input and generate prompt tokens, which are subsequently combined at the convolutional layers of the prompt networks $p_{a}$ and $p_{s}$, respectively.
In the event of transmitting the encoded RGB image to another device, the associated LiDAR map in sent as well. It is important to note that compressing a depth map typically results in a marginal increase in coding cost, even due to the considerably lower spatial resolution of the LiDAR map compared to the RGB image (see Sec. IV). A similar framework is discussed in [16], where indeed a semantic/instance label map of the original image is utilized as side information at both the encoder and decoder for extreme low rate coding.
@@
Note also that in the case of transmitting the LiDAR map alongside, an effective solution might involve jointly compressing both the RGB image and the LiDAR depth map. This approach would entail designing a model that takes both the RGB image and LiDAR map as inputs, encoding both of them, and producing both their reconstructed versions as outputs. Alternatively, one could envision a model that uses the LiDAR map exclusively at the encoder side, without involving it in the decoding process. In fact, despite the established principle derived from Berger’s result, indicating that side information, when solely available at the encoder and not influencing the distortion measure, does not offer benefits [17], one could explore whether incorporating LiDAR data at the encoder could improve the convergence of an end-to-end learning model and potentially result in a more efficient data compression model. We leave these potential investigations for future works.
III Proposed method
@@
III-A Background: reference architecture
@@
III-B LiDAR Depth Map Guided Image Compression Model
Fig. 2 depicts our image compression model guided by LiDAR depth maps. It draws inspiration from [15]. However, two significant distinctions set our approach apart from [15]. Firstly, our method does not incorporate the ROI mask, as it plays no role in our model, in contrast to the approach outlined in [15], which proposes a ROI-based coding. Secondly, the distinctive feature of our model lies in the inclusion of two supplementary prompt generation networks, referred to as $l_{a}$ and $l_{s}$ situated at the encoder and decoder sides, respectively. Both of them receive as input the LiDAR depth map $M_{D}\in\mathbb{R}^{1\times H\times W}$ and generate prompt tokens, which are then added to the prompt tokens of $p_{a}$ and $p_{s}$, respectively. These prompts allow the model to leverage the LiDAR information effectively for both the encoding and decoding processes.
@@ III-B LiDAR Depth Map Guided Image Compression Model
IV Experiments
IV-A Dataset
@@ IV-B Results
We conduct a single-stage training of our model using the ARKitScenes dataset’s training set. During each training iteration, input images are randomly cropped to a size of $256\times 256$ and the LiDAR maps undergo cropping in corresponding positions. Our entire model is trained jointly for variable rate coding by uniformly sampling $m_{\lambda}$ from 00 to $1$. To ensure fairness in comparison, we adhere to the same training procedure for the model without LiDAR prompts; that is, the model presented in [15] when considering uniform ROI maps.
Figure 3: Comparison of the model performance with and without LiDAR in terms of average RD curves, evaluating PSNR (a) and SSIM (b) as distortion metrics. Additionally, we include the performance of our model when provided with a uniform LiDAR map as side information for completeness. The results are evaluated on the $5,999$ images in the test set.
Figure 4: Visual results for two images from the test set are displayed in (a)-(b). The reconstructed images using LiDAR as side information are depicted in (c)-(d), while those reconstructed without LiDAR are shown in (e)-(f). The reconstructed images using a random map as side information are seen in (g)-(h). Corresponding LiDAR maps are represented in (i)-(j).
@@
Fig. 3 shows the rate-distortion (RD) curves, analyzing both PSNR-RGB (used in the training loss function) and SSIM, shown separately in Fig. 3a and Fig. 3b. The model using LiDAR consistently outperforms the model without LiDAR in both cases. In terms of Bjøntegaard’s metric, integrating LiDAR results in an average PSNR gain of $0.83$ dB and an average bit rate reduction of $16\%$ compared to its absence.
@@
Furthermore, we have evaluated the model’s performance by considering scenarios in which the original LiDAR data is unavailable at the decoder, such as during transmission. In such scenarios, the encoder typically employs the original LiDAR map as input to compress the RGB image. Next, the LiDAR map undergoes separate compression using our model, excluding the prompts $l_{a}$ and $l_{s}$, and employing $m_{\lambda}=0$ to minimize the rate incurred during LiDAR encoding. Then, it is transmitted to the decoder. At the decoder end, the received encoded LiDAR map is decoded and finally used as side information instead of the original LiDAR map. Interestingly, the resulting performance is comparable to that achieved by the model using the original LiDAR map, both in terms of PSNR and SSIM. In fact, the increase in coding cost while encoding the LiDAR depth map is only marginal, even due to the significantly lower spatial resolution of the LiDAR map in comparison to the RGB image. Moreover, using the decoded LiDAR map in place of the original one does not significantly affect the compression process. Notably, we remark that we have not retrained the model using the decoded LiDAR maps.

### 18 Mar 2024 v2

@@ Abstract
The incorporation of LiDAR technology into some high-end smartphones has unlocked numerous possibilities across various applications, including photography, image restoration, augmented reality, and more. In this paper, we introduce a novel direction that harnesses LiDAR depth maps to enhance the compression of the corresponding RGB camera images. To the best of our knowledge, this represents the initial exploration in this particular research direction. Specifically, we propose a Transformer-based learned image compression system capable of achieving variable-rate compression using a single model while utilizing the LiDAR depth map as supplementary information for both the encoding and decoding processes. Experimental results demonstrate that integrating LiDAR yields an average PSNR gain of $0.83$ dB and an average bitrate reduction of $16\%$ as compared to its absence.
Index Terms—  Learned image compression, LiDAR, depth map, prompts, transformer.
1 Introduction
@@
Fig. 1: High-level architectural framework. Our LiDAR Depth Map Guided Image Compression Model takes input from both the RGB image and the LiDAR depth data, using the LiDAR map to enhance the compression process of the RGB image. In the considered scenario, the LiDAR map is available at both the encoder and the decoder.
@@
The rest of the paper is structured as follows. Sec. 2 depicts the specific scenario under investigation. Sec. 3 introduces the architecture proposed for incorporating LiDAR information, and Sec. 4 shows the experimental results. Finally, Sec. 5 provides the concluding remarks.
2 Examined Scenario
Let us consider a simple scenario to assess the effectiveness of using a LiDAR map in compressing an RGB image, that is, we assume that the LiDAR map is available at both the encoder and the decoder, as illustrated in Fig. 1. In light of this scenario, we aim to investigate two distinct cases.
@@ 2 Examined Scenario
As a first case, we consider to have the LiDAR map available in its original uncompressed format. Thus, we examine the situation where the RGB image and the corresponding LiDAR map are simultaneously acquired from the device, and the uncompressed LiDAR map is stored without compression on the acquisition device, as it will be utilized for the various imaging manipulations previously mentioned. Then, the uncompressed LiDAR data will assist in both encoding the RGB image, which will be then locally stored on the device, and decoding the reconstructed version. In essence, this approach capitalizes on the presence of the LiDAR map on the device to enhance the compression of the RGB image. In this context, we will not consider the rate cost of the LiDAR map, as its presence on the device is independent of the RGB image compression goal.
@@
Fig. 2: The network architecture of the proposed LiDAR depth map guided image compression model. For variable rate control, the prompt generation networks $p_{a}$ and $p_{s}$ produce prompt tokens for the encoder $g_{a}$ and decoder $g_{s}$, respectively. Two additional prompt generation networks, denoted as $l_{a}$ and $l_{s}$, receive the LiDAR map as input and generate prompt tokens, which are subsequently combined at the convolutional layers of the prompt networks $p_{a}$ and $p_{s}$, respectively.
Secondly, we examine the case in which the RGB image must be transmitted to a second device. As transmitting the uncompressed LiDAR map solely for image replication on a remote device would be impractical due to the high transmission cost, we explore the impact of employing a compressed LiDAR map within the considered image compression framework and its associated rate cost. Thus, in this context, the LiDAR map is first compressed independently, and then the compressed version is used to assist the encoding on the first device and the decoding on the second device of the RGB image. It is important to highlight that, for the image encoding process, we would have the option to employ the uncompressed LiDAR map, assuming it has been stored on the acquisition device without compression (or at a very high quality). However, we opt to use the compressed LiDAR map for assisting the image encoding as well, ensuring uniformity in the LiDAR data used for both image encoding and decoding. It is also important to note that incorporating a compressed depth map into the bitstream typically results in a marginal increase in coding cost, even due to the considerably lower spatial resolution of the LiDAR map compared to the RGB image (see Sec. 4.1). A similar framework is discussed in [16], where indeed a semantic/instance label map of the original image is utilized as side information at both the encoder and decoder for extreme low rate coding.
@@
We would like to remark that in the case of transmitting the LiDAR map alongside, an effective solution might involve jointly compressing both the RGB image and the LiDAR depth map. This approach would entail designing a model that takes both the RGB image and LiDAR map as inputs, encoding both of them, and producing both their reconstructed versions as outputs. Alternatively, one could envision a model that uses the LiDAR map exclusively at the encoder side, without involving it in the decoding process. In fact, despite the established principle derived from Berger’s result, indicating that side information, when solely available at the encoder and not influencing the distortion measure, does not offer benefits [17], one could explore whether incorporating LiDAR data at the encoder could improve the convergence of an end-to-end learning model and potentially result in a more efficient data compression model. We leave these potential investigations for future works.
3 Proposed method
@@
3.1 Background: reference architecture
@@
3.2 LiDAR Depth Map Guided Image Compression Model
Fig. 2 depicts our image compression model guided by LiDAR depth maps. It draws inspiration from [15]. However, two significant distinctions set our approach apart from [15]. Firstly, our method does not incorporate the ROI mask, as it plays no role in our model, in contrast to the approach outlined in [15], which proposes a ROI-based coding. Secondly, the distinctive feature of our model lies in the inclusion of two supplementary prompt generation networks, referred to as $l_{a}$ and $l_{s}$ situated at the encoder and decoder sides, respectively. Both of them receive as input the LiDAR depth map $M_{D}\in\mathbb{R}^{1\times H\times W}$ and generate prompt tokens, which are then added to the prompt tokens of $p_{a}$ and $p_{s}$, respectively. Note that the LiDAR depth map $M_{D}$ undergoes an initial upsampling process to align its size with that of the RGB image. This is necessary because the resolution of LiDAR sensors is typically smaller ($256\times 192$ for Apple’s devices), compared to the RGB camera.
We would like to remark that in  [15], the ROI map information is concatenated with the lambda map $M_{\lambda}$ (only at the encoder side). However, we found that concatenating the LiDAR depth map with $M_{\lambda}$ does not allow the model to significantly exploit the depth information. Instead, incorporating the specific prompts $l_{s}$ and $l_{a}$ at both the encoder and decoder sides, which receive both $M_{D}$, enables the system to effectively leverage the LiDAR depth information (as we will show in the next section).
@@ 3.2 LiDAR Depth Map Guided Image Compression Model
Fig. 3: Comparison of the model performance with and without LiDAR in terms of average RD curves, evaluating PSNR (a) and SSIM (b) as distortion metrics. The blue curve corresponds to the model’s performance when given the uncompressed LiDAR map, while the red curve represents the model’s performance when supplied with the compressed LiDAR map, which includes the cost of transmitting LiDAR information. The results are evaluated on the $5,999$ images in the test set.
4 Experiments
4.1 Dataset
@@ 4.2 Results
We conduct a single-stage training of our model using the ARKitScenes dataset’s training set. During each training iteration, input images are randomly cropped to a size of $256\times 256$ and the uncompressed LiDAR maps undergo cropping in corresponding positions. Our entire model is trained jointly for variable rate coding by uniformly sampling $m_{\lambda}$ from 00 to $1$. To ensure fairness in comparison, we adhere to the same training procedure for the model without LiDAR prompts; that is, the model presented in [15] when considering uniform ROI maps.
@@
Fig. 3 shows the rate-distortion (RD) curves, analyzing both PSNR-RGB (used in the training loss function) and SSIM, shown separately in Fig. 3a and Fig. 3b. The blue curve represents the model adopting the uncompressed LiDAR for both the encoding and decoding stages. As it can be observed, this model consistently outperforms the model without LiDAR for both metrics. In terms of Bjøntegaard’s metric, integrating uncompressed LiDAR results in an average PSNR gain of $0.83$ dB and an average bit rate reduction of $16\%$ compared to its absence.
@@
Furthermore, the red curve corresponds to the performance of our model when considering the case in which the uncompressed LiDAR data is unavailable, such as during transmission. In this context, the LiDAR information needs to be compressed separately. In our experiments, it undergoes separate compression using our model, where we specifically provide zero matrices as input to the prompts $l_{a}$ and $l_{s}$ and set $m_{\lambda}=1$ to maximize the quality of the compressed LiDAR map. Then, the compressed LiDAR map is used for assisting the RGB image encoding and decoding. Of course, the final bitstream will consist of both the latent spaces of the RGB image and the LiDAR map. Interestingly, the resulting performance is comparable to that achieved by the model using the original LiDAR map, both in terms of PSNR and SSIM. In fact, the increase in coding cost while encoding the LiDAR depth map is only marginal (around 9% on average), even due to the significantly lower spatial resolution of the LiDAR map in comparison to the RGB image. Moreover, using the compressed LiDAR map in place of the original one does not significantly affect the compression process. Specifically, in terms of Bjøntegaard’s metric, integrating compressed LiDAR results in an average PSNR gain of $0.57$ dB and an average bit rate reduction of $10\%$ compared to its absence (see Tab. 1 for recap). Notably, we remark that we have not retrained the model using the compressed LiDAR maps.
Uncompressed LiDAR Compressed LiDAR
$\Delta$ PSNR [dB] 0.83 0.57
$\Delta$ rate [%] -16 -10
Table 1: BD rate associated to Fig. 3a. The baseline performance is set with the model that does not use the LiDAR map (no LiDAR).
Fig. 4: Visual results for two images from the test set are displayed in (a)-(b). The reconstructed images using LiDAR as side information are depicted in (c)-(d), while those reconstructed without LiDAR are shown in (e)-(f). The reconstructed images using a random map as side information are seen in (g)-(h). Corresponding LiDAR maps are represented in (i)-(j).